Introduction: Proton-Antiproton Scattering

The proton is a complex object

• Consists of valence quarks (uud), sea quarks and gluons • each type of parton has momentum distribution function • p beam is really “broad-band” beam of quarks and gluons • p̄ is primarily beam of anti-quarks and gluons

Interested mostly in “hard scattering”

• Characterised by high pT of interaction

Parton Distribution Functions

Fraction of momentum carried by parton given by pdf’s

• Example is EHLQ (Set I) at Q2 = 5 GeV2

• Derived from deep inelastic scattering

• Note large non-valence quark contribution – Largest uncertainty comes from gluon contribution

QCD predicts these PDF’s to evolve with Q^{2}

• Get “softer” as ŝ increases

• Best measurements come from neutrino and DIS

• Raison d’etres for several experiments

Evolution of PDF’s

QCD predicts that f_{q}(x, Q^{2}) will evolve as Q^{2} increases

• Valence quark distributions become softer

• Sea quark and gluon distributions increase in size

• Evolution predicted by the Altarelli-Parisi equations

As example, Duke-Owens I evolution

• Red is x u(x, Q2) and blue is x d(x, Q2)

• Plotted for Q2 = 4 and Q2 = 5000 GeV2